Find two consective odd integers whose product is 99
Answer by Alan3354(69443)

Find two consective odd integers whose product is 99
---------------
Method 1
Sqrt(99) = ~ 10
Integers = 9 and 11.
-------------------
Method 2
x*(x+2) = 99
x^2 + 2x - 99 = 0
(x-9)*(x+11) = 0
x = 9, -11
Integers = 9 and 11
Integers = -11 and -9